OpenStudy (anonymous):

can someone solve this one, factor the polynomial completely, y^4+y^3+y+1

OpenStudy (anonymous):

= ( y + 1) ( y³ + 1)

OpenStudy (anonymous):

can you explain how you got there please

OpenStudy (anonymous):

no forget it, you distributed

OpenStudy (anonymous):

@k3lvinf17 Question?

OpenStudy (anonymous):

solve the system of equations x-y+5z=4 3x+z=0 x+4y+z=-16

OpenStudy (anonymous):

As the rule, one question per post!

OpenStudy (anonymous):

oho ok